# LLM2TEA: Agentic AI Designer Finds Innovative Objects with Generative Evolutionary Multitasking

### v2

@@ III-B3 Selection
We draw inspiration from MFEA [38] and employ a scalar fitness measure to compare genotypes across domains. Specifically, given tasks and genotypes in the joint population where , a set of factorial ranks for the tasks is determined by finding the ordered position of every genotype with respect to all other genotypes’ factorial cost. Here, this factorial cost is the genotype’s fitness score. As such, each genotype in the joint population will have a set of factorial ranks for all tasks. Thereafter, the scalar fitness value for the -th genotype is computed as . The scalar fitness values of the joint population are used in any evolutionary selection technique to prune less fitting genotypes, allowing the surviving prompts to form the next generation population in the evolution process.
@@ IV-A Test Scenarios
The performance of LLM2TEA is observed in the context of engineering design instances. Following (1) and (2), the experiments are set up as a two-task design problem that aims to find innovative car and airplane designs concurrently in a single EMT run. As such, the car domain is , while the airplane domain will be . We demonstrate the effectiveness of LLM2TEA with ChatGPT-3.5 [39] to drive the evolution and Shape-E [14] as the auxiliary text-to-3D generative model. Hence, the search space for effective prompts is and . In terms of in Equation (1) and (2), BLIP-2 VLM [22] serves as the evaluator to assess the visual criteria of the arrived designs. This visual criterion has a contribution weight of , taking priority of the physics or geometric criteria in the objective. In addition, we set and in Equation (2) to constrain the exploration of generated objects within and near domain boundaries. Four test scenarios are studied to verify the effectiveness of the LLM2TEA:
Test Scenario I: This test scenario assess the diversity of designs of LLM2TEA in multi-objective setting that aims to minimize drag in both car and airplane designs, and maximize lift in airplane while minimize lift in car designs.
Test Scenario II: This test scenario considers minimization of the projected frontal area [40] as .
Test Scenario III: This test scenario minimizes drag aerodynamic criteria with the use of OpenFoam [41] as .
Test Scenario IV: This test scenario assesses the creativity capability of the LLM2TEA in the discovery of creative design. Specifically, we set the two-task optimization tasks as “banana car” and “banana airplane” in this test scenario. We ground the designs with real-world performance design optimization criteria to analyze the creativity of the designs while adhering to real-world requirements.
@@ IV-B Parameter Settings
In all experiments conducted, the population size is set to , and the same random seeds are used throughout the evolution with a terminating condition of up to 20 generations, resulting in up to 400 designs generated and evaluated from the text-to-3D generative model in a single run. On the other hand, the classic tournament selection algorithm [42] is employed as the environment selection technique to select genotypes that form the population for the next generation. Furthermore, in a single experiment run, the evolutions and simulations are performed in parallel on a single shared compute node with a configuration comprised of Intel Xeon Silver 64 CPU cores clocked at 2.10 GHz, 128GB of RAM, and 3 Nvidia Quadro GV100 GPUs (32 GB each).
@@ IV-C Evaluating Designs
In the context of LLM2TEA, an important observation is its ability to generate designs that deviate from its conventional form. This capability is particularly pronounced in a cross-domain knowledge-driven approach, where the integration of knowledge from multiple domains often results in designs with fusion features, leading to innovative outcomes. However, in existing generative design research in literature [43, 44, 45, 46, 47, 48, 49], there is no clear definition or standard for the concept of “innovative” or “novel” design. In this paper, we define designs by examining the distinction between “conventional” and “unconventional” shapes. Specifically, we consider shapes randomly sampled from a generative model based on a known domain t (e.g. A car or An airplane) as conventional, i.e., designs sampled as . Designs with shapes deviating from this conventional set are considered innovative.
To quantify innovation, we propose an innovation score based on this idea. Using a VLM, a generated 3D design is compared with the visual bias of a target concept specified in text form. Specifically, the VLM is tasked to measure the similarity distance between the design and the task label (such as “A car” or “A airplane”). This measurement can indicate whether the design is visually similar to a conventional concept of the vehicle task that VLM has already learned and whether the text-to-3D generative model can be synthesized. As such, an Innovation Score metric is introduced and formulated as:
where is the -th task label with the format “A ” and is the CLIP[21] VLM that measures the similarity distance between the design and the task label. This VLM is used in the Shape-E model to train and generate 3D designs, aligning with the visual bias inherited in the text-to-3D generative model. Notice when , the design is considered to exhibit conventional features only. Conversely, when , the design is considered to be “innovative” or “novel”, deviating from its unconventional form known to the VLM. On a separate note, the Monte Carlo estimation is adopted here to approximate using 1,000 generated samples per task as the mean baseline, taking into account the variety of conventional shapes that resemble the target task.
@@ V Results and Discussions
This section discusses the experiment results in the following order. First, we show the designs discovered in our experiments. Subsequently, we analyze the diversity of designs. Following that, we discuss the various components of the LLM2TEA to evaluate its capabilities and effectiveness. Finally, we showcase the experiments for 3D printing the generated designs.
@@ V-A Results of Discovered Designs
Some of the most representative examples from all test scenarios of vehicle designs uncovered in the text-to-3D generative model, Shape-E, are presented in Fig. 3. Vehicle designs from the baseline exhibit conventional vehicular features, as shown in Fig. 3(a) and (b), and have a distinct contrast with the innovative designs found with the LLM2TEA, as illustrated in Fig. 3(c) and (d). Such innovative designs deviating from its conventional form are the result of driving the generative models to synthesize complex prompts that lead to designs that not only satisfy the physical or geometric criteria but also the visual specification of a car or airplane. Notice the difference in designs generated with LLM2TEA and the baseline. The cross-domain knowledge learning capability of LLM2TEA enables the genetic material of parent prompts from different domains to undergo crossover and mutation. This process generates offspring prompts that drive the discovery of hybrid vehicle designs. These designs feature unconventional body shapes and structures, such as banana-like curvatures or extended rear spoilers, challenging traditional car and airplane design conceptions.
Remark: In LLM2TEA, the implicit genetic transfer occurring during cross-domain mating facilitates the creation of hybrid designs that combine the advantageous traits of each parent genotype from different domains. On a separate note, LLM and text-to-X generative models known for its hallucination phenomenon [50] may lead to synthesizing outputs that deviate from the factual or the input prompt. Such a hallucination outcome is not always detrimental but can foster positive creativity, as highlighted by Jiang et al. [51]. Both of these algorithmic properties create the opportunity for emerging design trends, facilitating the discovery of innovative designs.
@@
V-B Diversity of Designs
In Test Scenario I, we compare the last five generations of best-performing designs on the multi-objective setting in a single run discovered using LLM2TEA against the baseline designs generated from a vanilla text-to-3D generative model without LLM2TEA. As shown in Fig. 4, the designs discovered with the LLM2TEA possess better aerodynamic performance than the baseline designs. Specifically, the airplane designs have demonstrated higher normalized lift () and lower normalized drag () than baseline designs. The same observations are made in the car designs. Apart from that, notice that the vanilla text-to-3D generative model discovers most of the designs in localized regions. In contrast, the LLM2TEA allows more diverse designs, as observed in the visual inspection of the generated airplane designs. We further examine the diversity performance of the LLM2TEA with the baseline designs. In Table I, we present the hypervolume indicator [52] measured against the physics objectives ( and ) space, which ranges from 0.0 to 1.0, and discovers and times more hypervolume values for airplane and car designs, respectively. This demonstrated that our proposed LLM2TEA effectively explores design spaces to uncover innovative designs with good physics indicators.
V-C Effectiveness of LLM2TEA
A series of experiments examines various LLM2TEA capabilities in driving GenAI models to discover innovative solutions objectively by comparing a cross-domain knowledge-driven approach with a baseline. The findings are presented here.
V-C1 Effect of Evaluation Criteria
Here, we examine the effect of the objective functions defined in Section III-A, showing that these objectives are able to guide the LLM2TEA to find the best-performing designs. As demonstrated in Fig. 5, with visual criterion only, the LLM2TEA will discover vehicle designs that have wide body shapes and high roofs that yield poor aerodynamic performance while still retaining the characteristics of a car or airplane (shown in Fig. 5(a)). In contrast, the vehicle designs evaluated with physical criterion only (shown in Fig. 5(b)) are impractical. Hence, considering both the physical and visual criteria will lead the LLM2TEA to uncover practical designs with body shapes that yield the lowest drag performance (shown in Fig. 5(d)) as compared to the designs in Fig. 5(a) to (c). As a matter of fact, more than 73% of the generated vehicle designs have better drag performance than the top 1% percentile of the designs in the baseline. This result suggests that the LLM2TEA effectively uncovers innovative yet practical designs.
V-C2 Evolution Trends of LLM2TEA
The LLM2TEA guidance to GenAI models in yielding fitter designs can be observed from the average fitness scores across generations in the single run. As presented in Table II, a distinct trend of gradual improvement in the initial phases, particularly for the airplane design task. This progression becomes more pronounced in the latter half of the run, indicating that the evolution process led to the discovery of better aerodynamic designs.
While the LLM2TEA instructs the LLM to construct the prompts with respect to the objective, the initial samples generated by the text-to-3D generative model (Shape-E) in all our test scenarios (Fig. 6, 7 and 8) still resulted in vehicle designs with conventional features. We observe that the prompts evolve with LLM2TEA in the evolutionary run and notice the introduction of disruptive words in these prompts, as early in the generations, could have led the designs to deviate from its conventional form. For instance, in Fig. 6, Shape-E is initially tasked to generate a dart-like shape airplane. However, the initial design resembles a Concorde jet more than a dart-shaped airplane. After a few generations in the early phase, the airplane wing design took the blade-form shape and a pointed nose, resulting in a smaller projected frontal area body shape. When serving as parents for future generations, such a design led to the consistent use of words that favored extremely sleek body shapes with sharp and pointed noses, as such words tend to represent minimal geometric shapes. This evolutionary discovery approach applied to other test scenarios as presented in Fig. 7 and 8 reproduced similar design transformation behavior.
@@ V-C3 Behavior of Mating Operations
In Fig. 7, notice the less frequently used words included in the car prompts, such as ‘wings’ and ‘jet’. Such prompt constructs are the result of the LLM2TEA mating operations. Self-mating and same-domain mating operations construct same-domain prompts that describe using commonly used terms, resulting in vehicle designs with body shapes that tend to look more conventional. In contrast, cross-domain mating operation, with its cross-concept knowledge integration, allows the construction of prompts by transferring representative words from the primary task to the other secondary task. The resulting designs exhibit unconventional body shapes or structures that retain the visual look of the primary task and satisfy the physics or geometric criteria. As illustrated in Fig. 9 of the cross-domain mating operation, the LLM is instructed to select words in both parents in the crossover operation randomly. The LLM tends to generate a prompt that can yield a fitter design while retaining the parents’ genetic properties. This demonstrated the LLM’s capability in performing proper cross-concept knowledge integration during cross-domain mating, resulting in a 3D design with the hybrid design look of the parents.
Since the LLM2TEA cross-concept knowledge transfer mechanism operates by transferring representative words from one task to another, the effectiveness of this capability can be observed by examining the percentage overlap between each task’s vocabulary set. The vocabulary set for each task comprises representative word tokens of all the prompts generated throughout the evolutionary run. For a fair comparison, frequently occurring stopwords (such as ‘a’, ‘the’, ‘is’, etc) are pruned from the vocabulary set. Furthermore, all word tokens are reduced to its base word form by transforming the morphological variants to the base word. The resultant vocabulary set comprises the word tokens most representative of each task. Henceforth, we observe 38.38% overlap in the vocabulary sets between tasks, indicating a proportion of word tokens are shared between tasks. The result demonstrated the effectiveness of the cross-domain knowledge-driven approach, which triggers a significant amount of words to be transferred between tasks with LLM2TEA.
| (a) | (b) | (c) |
@@ V-C4 Efficiency and Scalability of LLM2TEA
From our observations of the computational time of each key component in the LLM2TEA framework, the operators’ computation barely took up any time as we utilize proprietary LLMs, which run on third-party servers. However, 72.7% of one generation total time is used in the OpenFoam physics simulation, which is typical in engineering design optimization problems. The remaining time is used in visual evaluations, taking up 20.5% of the time, and text-to-3D generation utilizes the remaining 6.8% of the time. Similar to existing engineering design methods, where scaling up for bigger problems requires parallelizing physical simulation computations, LLM2TEA efficiently uses multicore CPU processing to compute the physical performance of generated objects. However, in contrast to existing evolutionary methods that require a large population size to yield performance, LLM2TEA is able to achieve good results in small population sizes in the tens of orders. In addition, as the generated objects are available in digital 3D formats, we can scale the objects down to fit into a unit cube, further reducing the simulation computation requirements.
V-C5 Quantization of Innovative Solutions
The performance of discovering innovative solutions with LLM2TEA is examined from a visual perspective. Using the proposed Innovation Score metric in Equation (3), we observe more than 69.12% and 46.23% of airplanes and cars found with the LLM2TEA to be above this mean baseline, respectively. The results underscore the effectiveness of the evolutionary process using LLM2TEA in guiding generative models to discover cross-domain innovative designs. This also provided the justification for the effectiveness of the LLM2TEA that leads to innovative designs being discovered within same-domain or cross-domains, such as the ones discussed in Fig. 3(c) and (d) with unconventional features, such as more streamlined, futuristic, and sleeker body shapes.

### v3

@@ III-B3 Selection
A scalar fitness measure to compare genotypes across domains is considered. Specifically, given tasks and genotypes in the joint population where , a set of factorial ranks for the tasks is determined by finding the ordered position of every genotype with respect to all other genotypes’ factorial cost. Here, this factorial cost is the genotype’s fitness score. As such, each genotype in the joint population will have a set of factorial ranks for all tasks. Thereafter, the scalar fitness value for the -th genotype is computed as . The scalar fitness values of the joint population are used in any evolutionary selection technique to eliminate less fitting genotypes, allowing the surviving individuals to form the next generation of population in the evolutionary process.
@@ IV-A Test Scenarios
The performance of LLM2TEA is observed in the context of engineering design instances. Following Equations (1) and (2), the experiments are set up as a two-task design problem that targets to find novel car and airplane designs concurrently in a single EMT run. As such, the car domain is , while the airplane domain is . The effectiveness of LLM2TEA with ChatGPT 3.5 [36] in driving evolution and Shape-E [14] as the auxiliary text-to-3D generative model is considered. The search space for effective prompts is thus and for Shape-E, while a maximum sequence context length of and for ChatGPT 3.5. In terms of in Equation (1) and (2), BLIP-2 VLM [22] is used as the evaluator to assess the visual performance of the generated designs. This visual criterion has a contribution weight of , taking priority over the physical criteria in the objective. In addition, and in Equation (2) are set to constrain the exploration of generated designs within and near domain boundaries. Four test scenarios are considered to study the effectiveness of the LLM2TEA:
Test Scenario I: This test scenario assesses the design diversity attained by LLM2TEA in a multi-objective setting that aims to minimize aerodynamic drag in both car and airplane designs, simultaneously maximizing aerodynamic lift in airplanes while minimizing aerodynamic lift in car designs.
Test Scenario II: This test scenario considers a minimization of the projected frontal area [37] as .
Test Scenario III: This test scenario minimizes aerodynamic drag criterion.
Test Scenario IV: This test scenario assesses the creativity capability of LLM2TEA in discovering novel designs. Specifically, the two-task optimization tasks are defined as “banana car” and “banana airplane”. The generated designs are grounded in real-world performance-based optimization criteria, allowing for an assessment of creativity while ensuring conformity to real-world requirements.
Aerodynamic drag and lift of generated designs are computed using OpenFOAM [38] and normalized. OpenFOAM is an open-source computational fluid dynamics software toolkit primarily used for simulating fluid flow, heat transfer, turbulence, chemical reactions, and other related physical phenomena.
@@ IV-B Parameter Settings
In all experiments, the population size is set to , and consistent random seeds are used throughout the evolution, with a terminating criterion of up to generations. This results in up to designs being generated and evaluated by the text-to-3D generative model in a single run. The classic tournament selection algorithm [39] is used to select the genotypes that form the next generation of designs. Each experiment run is executed on a single shared compute node equipped with Intel Xeon Silver 64 CPU cores clocked at 2.10 GHz, 128 GB of RAM, and three Nvidia Quadro GV100 GPUs (each with 32 GB of RAM), with both evolution and simulation run in parallel.
@@ IV-C Evaluating Designs
In the context of LLM2TEA, an important observation is its ability to generate designs that deviate from its conventional form. This capability is particularly evident in cross-domain design scenarios, where integrating functional principles and constraints from multiple domains often produces artifacts with fused structural or behavioral traits, resulting in novelty. However, in existing generative design research in literature [40, 41, 42, 43, 44, 45, 46], there is no clear definition or standard for the concept of “novel” design. This paper defines design novelty by examining the distinction between “conventional” and “unconventional” shapes. Specifically, randomly sampling shapes from a generative model based on a known domain t (e.g., a car or an airplane) is considered conventional, i.e., designs sampled as . Designs with shapes deviating from this conventional set are considered novel designs.
To quantify novelty, a Novelty Score metric is proposed based on this idea. Using a VLM, a generated 3D design is compared with the visual bias of a target concept specified in text form. Specifically, the VLM is tasked to measure the similarity distance between the design and the task label (such as “A car” or “A airplane”). This measurement can indicate whether the design is visually similar to a conventional concept of the vehicle task that VLM has already learned and whether the text-to-3D generative model can be synthesized. As such, Novelty Score metric is introduced which is formulated as:
where is the -th task label with the format “A ” and is the independent VLM (e.g., CLIP[21]) to measure the similarity distance between a design and its corresponding task label. This VLM is used in the Shape-E model to train and generate 3D designs, aligning with the visual bias inherent in the text-to-3D generative model. Notice that when , the design is considered to exhibit only conventional features. Conversely, when , the design is considered to be “novel”, indicating a deviation from the conventional forms known to the VLM. On a separate note, Monte Carlo estimation is adopted here to approximate using generated samples per task as the baseline mean, taking into account the variety of conventional shapes that resemble the target task.
@@ V Results and Discussions
This section presents the experiment results in the following order. First, the designs discovered during the experiments are showcased. Next, the diversity of these designs is analyzed, followed by a discussion of the various components of the LLM2TEA to evaluate its capabilities and effectiveness. Finally, experiments demonstrating the 3D printing of the generated designs are presented.
@@ V-A Results of Discovered Designs
Some of the most representative examples from all test scenarios of vehicle designs uncovered in the text-to-3D generative model, Shape-E, are presented in Fig. 3. Vehicle designs from the baseline exhibit conventional vehicular features, as shown in Fig. 3(a) and (b), and have a distinct contrast with the novel designs found with the LLM2TEA, as illustrated in Fig. 3(c) and (d). Such novel designs that deviate from conventional form result from guiding the generative models to synthesize complex prompts. These prompts lead to designs that not only satisfy the physical criterion but also fulfill the visual specification of a car or an airplane. A clear distinction can be observed between designs generated with LLM2TEA and those from the baseline model. The cross-domain knowledge learning capability of LLM2TEA enables the genetic material from parent prompts across different domains to undergo crossover and mutation, thereby facilitating the transfer of knowledge across domains. This process generates offspring prompts that drive the discovery of hybrid vehicle designs. These designs feature unconventional body shapes and structures, such as banana-like curvatures or extended rear spoilers, challenging traditional car and airplane design concepts.
Remark: In LLM2TEA, the implicit genetic transfer occurring during cross-domain mating facilitates the creation of hybrid designs that combine the distinctive traits of each parent genotype from different domains. On a separate note, LLM and text-to-X generative models known for its hallucination phenomenon [47] may lead to synthesizing outputs that deviate from the factual or the input prompt. Such a hallucination outcome is not always detrimental but can foster positive creativity [48]. Both of these algorithmic properties create the opportunity for emerging design trends, facilitating the discovery of novel designs.
V-B Diversity of Designs
In Test Scenario I, the best-performing designs from the last five generations in a multi-objective setting of a single run using LLM2TEA are compared against baseline designs generated with a text-to-3D generative model without LLM2TEA. As shown in Fig. 4, the designs discovered with the LLM2TEA exhibit superior aerodynamic performance compared to the baseline. Specifically, airplane designs demonstrated better lift () and drag () than baseline designs, with similar trends observed for car designs. Apart from that, the baseline text-to-3D generative model tends to discover designs clustered in localized regions, whereas LLM2TEA produces more diverse designs, as evident from the visual inspection of the generated airplane designs. To quantitatively assess the diversity, Table I presents the hypervolume indicator [49] measured over the physics objectives ( and ) space, which ranges from to . LLM2TEA achieves hypervolume values greater for airplane designs and times for car designs compared to the baseline. These results demonstrated that the proposed LLM2TEA effectively explores design spaces to uncover novel designs with strong physics performance.
V-C Effectiveness of LLM2TEA
A series of experiments was conducted to objectively examine the capabilities of LLM2TEA to discover novel solutions. The findings are presented here.
@@ V-C Effectiveness of LLM2TEA
V-C1 Effect of Evaluation Criteria
This subsection examines the effects of the objective functions defined in Section III-A, demonstrating their ability to guide LLM2TEA toward finding the best-performing designs. As shown in Fig. 5, when guided solely by visual criteria, LLM2TEA generates vehicle designs with wide body shapes and high roofs that retain the characteristics of a car or airplane but exhibit poor aerodynamic performance (Fig. 5(a)). In contrast, designs evaluated solely on physical criteria will result in the production of more highly fragmented designs, such as the example presented in Fig. 5(b). Therefore, incorporating both physical and visual criteria enables LLM2TEA to discover practical designs with body shapes that achieve the lowest drag (Fig. 5(d)) compared to those in Fig. 5(a) through (c). Notably, over of the generated vehicle designs outperform the top designs produced by the text-to-3D baseline in terms of physical performance. These results suggest that LLM2TEA effectively uncovers novel yet practical designs.
V-C2 Evolution Trends of LLM2TEA
Table II presents LLM2TEA average fitness scores across generations in the single evolutionary run. A distinct trend of gradual improvements in the initial phases can be observed, particularly for the airplane design task. This progression becomes more pronounced in the latter half of the run, indicating that the evolutionary search led to the discovery of better aerodynamic designs. Further examination of the results in all test scenarios supports the observed trend (Figs. 6, 7, and 8). While the initial vehicle designs exhibit body shapes resembling conventional forms, LLM2TEA enables the designs to evolve and deviate from their conventional forms, with more significant changes observed from the mid-generation phase of the evolutionary search.
@@ V-C2 Evolution Trends of LLM2TEA
| (a) | (b) | (c) |
@@ V-C3 Behavior of Mating Operations
In Fig. 7, notice the less frequently used words included in the car prompts, such as “wings” and “jet”. Such prompt constructs are the result of the LLM2TEA mating operations. Self-mating and same-domain mating operations construct same-domain prompts that utilize commonly used terms, resulting in vehicle designs with conventional body shapes. In contrast, cross-domain mating operation, with its cross-concept knowledge integration, enables the crossbreeding between different domains by combining shape-defining words from these domains to produce a new genotype. The resulting phenotype generated with this new genotype exhibits unconventional body shapes with a hybrid appearance, combining distinctive features of these domains while still satisfying the physical criteria (Fig. 9).
Since the LLM2TEA cross-concept knowledge transfer mechanism operates by transferring representative words from one task to another, the effectiveness of this capability can be observed by examining the percentage overlap between each task’s vocabulary set. The vocabulary set for each task comprises representative word tokens of all the prompts generated throughout the evolutionary run. For a fair comparison, frequently occurring stopwords (such as “a”, “the”, “is”, etc) are pruned from the vocabulary set. Furthermore, all word tokens are reduced to its base word form by transforming the morphological variants to the base word. The resultant vocabulary set comprises the word tokens most representative of each task. Henceforth, an overlap in the vocabulary sets between tasks, indicating a proportion of word tokens shared between tasks. The results demonstrated the effectiveness of the cross-domain knowledge-driven approach, which transfers a significant amount of words between tasks with LLM2TEA.
@@ V-C4 Efficiency and Scalability of LLM2TEA
From our observations of the computational time of each key component in LLM2TEA, the computation time of the operators is negligible as proprietary LLMs is used, which runs on third-party servers. However, of one generation total time is used in the OpenFoam physics simulation, which is typical in engineering design optimization problems. The remaining time is used for visual evaluations, accounting for of the time, and text-to-3D generation utilizes the remaining of the time. Similar to existing engineering design methods, where scaling up for bigger problems requires parallelizing physical simulation computations, LLM2TEA efficiently uses multicore CPU processing to compute the physical performance of generated designs. However, in contrast to existing evolutionary methods that require a large population size to yield good results, LLM2TEA can achieve good results with a small population size. Additionally, since the generated designs are available in digital 3D formats, the designs can be scaled down to fit into a unit cube, further reducing the simulation computation requirements.
V-C5 Quantification of Novel Solutions
The effectiveness of LLM2TEA in discovering novel solutions is examined from a visual perspective using the proposed Innovation Score metric defined in Equation (3). Results indicate that of airplane designs and of car designs generated by LLM2TEA exceed the computed mean baseline, underscoring the efficacy of the multitask evolutionary process in guiding generative models toward the discovery of cross-domain novel designs. This further validates LLM2TEA’s capability to yield unconventional solutions not only across domains but also within a single domain, as illustrated in Fig. 3(c) and (d), where the resulting designs exhibit more streamlined, futuristic, and sleeker body shapes that deviate from conventional norms.